OpenStudy (anonymous):

HELP! I though I got it right but it isn't a choice! which factor would you cancel from the numerator and denominator to simplify x^2-3x-4 over x^2-16

OpenStudy (mertsj):

Did you factor the numerator?

OpenStudy (anonymous):

I had thought it was x^2

OpenStudy (anonymous):

but it isn't a chioce

OpenStudy (mertsj):

Do you know what factors are?

OpenStudy (anonymous):

ya...

OpenStudy (mertsj):

So how can you think that x^2 is the factors of the numerator?

OpenStudy (mertsj):

If I tell you that the factors of 12 are 10 and 7, what would you say?

OpenStudy (anonymous):

cause its the same on the num. and the dem

OpenStudy (anonymous):

that isn't correct

OpenStudy (mertsj):

Exactly, because 10 x 7 is not 12

OpenStudy (mertsj):

So when you write factors, they must multiply back to give the original expression. So your first job in this problem is to find the factors of the numerator and then find the factors of the denominator.

OpenStudy (mertsj):

x^2 is not a factor. Factors are things that you multiply.

OpenStudy (anonymous):

so for 3 would factors be 1? and oohh I see

OpenStudy (mertsj):

\[x^2-3x-4=(x-4)(x+1)\]

OpenStudy (mertsj):

\[x^2-16=(x-4)(x+4)\]

OpenStudy (anonymous):

so would it be x-4?

OpenStudy (anonymous):

for the anwser

OpenStudy (mertsj):

Yes. And you had better learn to factor or you cannot go much further in math.

OpenStudy (anonymous):

4 sqrt 13+ sqrt 6-7 sqrt 13.. could you help me with this?

OpenStudy (mertsj):

Let me see if I know what the problem is:

OpenStudy (anonymous):

and yes I know...I do know how to factor I have a 97 in the FLVS class I was just having a brain fart.. sorry for making you do all this work but u are helping me

OpenStudy (mertsj):

\[4\sqrt{13}+\sqrt{6}-7\sqrt{13}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes that's right equation! :)

OpenStudy (mertsj):

It is not an equation since there is no equal sign. It is an expression. Do you know what similar radicals are?

OpenStudy (anonymous):

haha ya ik! and yes I do... I learned that in module 10 9.01 in FLVS

OpenStudy (anonymous):

or maybe 9.02 not sure

OpenStudy (mertsj):

Only similar radicals can be added and you add them by adding the coefficients.

OpenStudy (anonymous):

ok so would the answer be -3 sqrt 6?

OpenStudy (mertsj):

The two radicals that are similar are the first one and the last one.

OpenStudy (anonymous):

or wait -3 sqrt 13 + sqrt 6

OpenStudy (mertsj):

\[4\sqrt{13}+\sqrt{6}-7\sqrt{13}=4\sqrt{13}-7\sqrt{13}+\sqrt{6}=\sqrt{13}(4-7)+\sqrt{6}=-3\sqrt{13}+\sqrt{6}\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so the answer -3 sqrt 13 + sqrt 6

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ok! thank you very much u have no idea how much u helped me!

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

on: 4x-12 over 3x-9

OpenStudy (anonymous):

I think the answer is 8 over 3

OpenStudy (anonymous):

but I think im wrong

OpenStudy (anonymous):

Mertsj ?

OpenStudy (anonymous):

help!

OpenStudy (mertsj):

Factor 4 from the numerator and factor 3 from the denominator